9. Hedging strategy to protect against falling prices Price fluctuations in commodities can have significant consequences for companies, especially if the fluctuation involves a prime raw material for a company. Different companies will adopt different strategies to manage the risk in price fluctuations, indluding adjusting the timing of their commodity purchases, maintaining a safety stock of their raw materials, and hedging Consider the case of Cranked Coffee Company, a large copper-producing company The companys cost of producing copper is about $3.00 per pound. The current market price for copper is $3.60 per pound. The six-month futures price for copper is $3.75 per pound. At this selling price, the company can maintain its earnings growth. The company expects to produce 1,250,000 pounds of copper in this six months. (Note: Copper futures are traded at a standard size of 250,000 pounds.) If the company does not hedge the copper it produces, it can expect to eam a total revenue of end of six months. at the If Cranked Coffee places a hedge on its copper production in the futures market, it would delivery in six months at a delivery price of $3.75 per pound to generate profits that maintain its desired eamings growth. contracts for When the contract comes due in six months, the spot price of copper is $2.55 per pound in the cash markets. Prices on the new six-month futures contracts in copper are $3.19 per pound. Calculate the expected revenue in the following markets Cash Market Futures Market Net gain or loss in the cash market O $375,000 O $700,000 O-$562,500 O-$3,987,500 Net gain or loss in the futures market: O -$4,687,500 O $1,500,000 o $3,987,500 O -$562,500 The cost of production of copper is $3,750,000. Thus, Cranked Coffee will in the futures market and in the cash market. This gain and loss offset each other, and the company benefits from placing the hedge. This hedging strategy would be referred to as a hedge, and it helps protect the producer to sell a commodity against falling prices.
Answer #1

1. If company does not hedge it can earn total revenue of $4,500,000 (1250000*3.60)

2. To hedge the company will sell 5 (1250000/250000) contracts.

3. Loss in cash market = 1250000*(2.55-3)= -$562,500

Profit in future market = 5*(3.75-2.55)*250000= $1,500,000

4. Cranked coffee will sell in the futures market and do nothing in the cash market as it is already producing the copper i.e. it already has spot buy position in copper.

5. This hedge is referred to as "Short Hedge".